OpenStudy (anonymous):

Which of the following values "completes the square," or creates a perfect square trinomial, for x2 − 8x + ___? 8 16 –8 –16

OpenStudy (anonymous):

For x² + bx + c, the number that completes the square is \[\left( \frac{ b }{ 2 } \right)^2\] for your problem that's\[\left( \frac{ -8 }{ 2 } \right)^2\] for your problem that's

OpenStudy (anonymous):

so c is correct

OpenStudy (anonymous):

imeant b

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

yes 16

OpenStudy (anonymous):

can you answer another one

OpenStudy (anonymous):

Using the completing-the-square method, find the vertex of the function f(x) = 2x2 − 8x + 6 and indicate whether it is a minimum or a maximum and at what point. Maximum at (2, –2) Minimum at (2, –2) Maximum at (2, 6) Minimum at (2, 6)

OpenStudy (anonymous):

is it a

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

can't be a maximum. It opens up because the number in front of x² is positve

OpenStudy (anonymous):

so c would be correctr

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

can't be a maximum

OpenStudy (anonymous):

so that only leave b :) @peachpi

OpenStudy (anonymous):

yes